OpenStudy (anonymous):

∫ sin^6x cos^3x dx =∫sin^6x (1-sin^2x) cosx dx let,sinx=z =>d/dx sinx=dz/dx =>cosx dx=dz so,∫ sin^6x cos^3x dx=∫sin^6x cosx dx-∫sin^6x sin^2x cosx dx =∫z^6 dz-∫z^8 dz =z^6+1/6+1-z^8+1/8+1 +c =z^7/7+z^9/9+c =sin^7x/7+sin^9x/9 +c is this right?

hartnn (hartnn):

yes! thats absolutely correct :) good work!

hartnn (hartnn):

just that the sign in between the term should be minus

OpenStudy (anonymous):

Thank you very much :) genius sir :)

hartnn (hartnn):

=∫z^6 dz-∫z^8 dz =z^6+1/6+1-z^8+1/8+1 +c =z^7/7-z^9/9+c =sin^7x/7-sin^9x/9 +c

OpenStudy (anonymous):

oh sorry my head is hot little :)

hartnn (hartnn):

no problem :) welcome ^_^

OpenStudy (anonymous):

^_^ thank you again